OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)? A) -x + y = 2 B) -x – y = 2 C) x – y = 2 D) -x – y = 0 E) x + y = 0

OpenStudy (imstuck):

The first thing you need to do here is find the slope of the line you have. To do this you need to get it into slope-intercept form: y = mx + b, where m is the slope and b is the y-intercept. Doing this you get y = x + 7. What is the slope of the line y = x + 7?

OpenStudy (anonymous):

0?

OpenStudy (texaschic101):

if you are just stuck with a variable. like x, it is not written but it is there...it is 1x. x is the same as 1x...y is the same as 1y. So your slope of y = x + 7...which is the same as y = 1x + 7...making your slope 1 understand ?

OpenStudy (anonymous):

yes, thank you! I just don't undersand where to go from here!

OpenStudy (texaschic101):

now that we have the slope, we will continue. A perpendicular line will have a negative reciprocal slope. All that means is " flip " the slope, and change the sign. Slope = 1, or 1/1...so we flip the slope and change the sign, making our slope be -1. Have I lost you ?

OpenStudy (anonymous):

nope, I still understand!

OpenStudy (texaschic101):

now we will use y = mx + b slope(m) = -1 (-1,-1)..x = -1 and y = -1 now we sub -1 = -1(-1) + b -1 = 1 + b -1 - 1 = b -2 = b so our perpendicular line is : y = -x - 2...but it seems we need it in standard form Ax + By = C y = -x - 2 --- add x to both sides x + y = -2 hmmm..not an answer choice x + y = -2 --->(-1) -x - y = 2...well, that works

OpenStudy (anonymous):

thank you so soo much!!!

OpenStudy (texaschic101):

no problem :)